# Zero forcing irredundant sets

### v1

@@ 1 Introduction
A set $S$ is a dominating set of a graph $G$ if every vertex of $G$ is in $S$ or a neighbor of a vertex in $S$. A set $S$ of vertices in a graph G is irredundant if each vertex in $S$ dominates a vertex of G that is not dominated by any other vertex in $S$ (this may be a neighbor or itself), called a private neighbor. The study of irredundance was introduced in 1978 by Cockayne, Hedetniemi and Miller [11] in relation to the study of minimal dominating sets. The upper and lower irredundance numbers are the maximum cardinality and minimum cardinality of a maximal irredundant set over all maximal irredundant sets. Zero forcing was introduced as a bound for the maximum nullity among real symmetric matrices whose pattern of nonzero off-diagonal entries is described by the edges of a graph in [2] and independently in other applications. We introduce zero forcing irredundant sets with a private fort playing the role of a private neighbor; the upper and lower zero forcing irredundance numbers are the maximum cardinality and minimum cardinality of a maximal zero forcing irredundant set over all maximal zero forcing irredundant sets. A graph $G=(V(G),E(G))$ is simple, undirected, finite, and $V(G)\neq\emptyset$.
@@
Vertices $v$ and $u$ of $G$ are neighbors if $vu$ is an edge of $G$, the (open) neighborhood of $v$ is $N_{G}(v)=\{u\in V(G):u\mbox{ and }v\mbox{ are neighbors}\}$, the closed neighborhood of $v$ is $N_{G}[v]=N_{G}(v)\cup\{v\}$, and $N_{G}[S]=\cup_{x\in S}N_{G}[x]$ for $S\subseteq V(G)$. When the graph $G$ is clear, open and closed neighborhoods can be denoted by $N(v),N[v],N[S]$.
@@ 4 Bounds
First, we assume that $c_{j-1}$ and $a_{j}$ cannot be in a private fort of the vertices in $V_{j}\cap S$, and show $|V_{j}\cap S|\leq 1$. Let $x\in\{b_{j},c_{j},d_{j}\}\cap S$ and let $F_{x}$ be a private fort of $x$. If $F_{x}\cap\{b_{j},d_{j}\}\neq\varnothing$, then $a_{j}$ must be adjacent to two vertices in the fort and $c_{j-1}\not\in F_{x}$, so $\{b_{j},d_{j}\}\subseteq F_{x}$. If $x=c_{j}$, then $b_{j}\in F_{x}$ or $b_{j}$ needs two neighbors in $F_{x}$ so $d_{j}\in F_{x}$, so $\{b_{j},c_{j},d_{j}\}\subseteq F_{x}$. Therefore $|V_{j}\cap S|\leq 1$. By symmetry, if $c_{j-1}$ and $a_{j}$ cannot be in a private fort of the vertices in $V_{j-1}\cap S$, then $|S\cap V_{j-1}|\leq 1$. Thus {linenomath}

### v2

@@ 1 Introduction
A set $S$ is a dominating set of a graph $G$ if every vertex of $G$ is in $S$ or a neighbor of a vertex in $S$. A set $S$ of vertices in a graph G is irredundant if each vertex in $S$ dominates a vertex of G that is not dominated by any other vertex in $S$ (this may be a neighbor or itself), called a private neighbor. The study of irredundance was introduced in 1978 by Cockayne, Hedetniemi and Miller [11] in relation to the study of minimal dominating sets. The upper and lower irredundance numbers are the maximum cardinality and minimum cardinality of a maximal irredundant set over all maximal irredundant sets. Zero forcing was introduced as a bound for the maximum nullity among real symmetric matrices whose pattern of nonzero off-diagonal entries is described by the edges of a graph in [2] and independently in other applications. We introduce zero forcing irredundant sets with a private fort playing the role of a private neighbor. The upper and lower zero forcing irredundance numbers are the maximum cardinality and minimum cardinality of a maximal zero forcing irredundant set over all maximal zero forcing irredundant sets. A graph $G=(V(G),E(G))$ is simple, undirected, finite, and $V(G)\neq\emptyset$.
@@
It is immediate that $D\subseteq V(G)$ is a dominating set if and only if $D$ intersects every closed neighborhood.
A minimal fort is a fort that is not properly contained in any other fort. Since every fort contains a minimal fort, the next result is an immediate consequence of the previous theorem.
Corollary 1.2.
Let $G$ be a graph. Then $B\subseteq V(G)$ is a zero forcing set if and only if $B$ intersects every minimal fort.
@@
Vertices $v$ and $u$ of $G$ are neighbors if $vu$ is an edge of $G$, the (open) neighborhood of $v$ is $N_{G}(v)=\{u\in V(G):u\mbox{ and }v\mbox{ are neighbors}\}$, the closed neighborhood of $v$ is $N_{G}[v]=N_{G}(v)\cup\{v\}$, and $N_{G}[S]=\cup_{x\in S}N_{G}[x]$ for $S\subseteq V(G)$. When the graph $G$ is clear, open and closed neighborhoods can be denoted by $N(v),N[v],N[S]$. Note that $S$ is irredundant if and only if for every $u\in S$ there exists a $v\in V(G)$ such that $S\cap N[v]=\{u\}$, so the closed neighborhood of a private neighbor can also be seen as a model for a private fort in the context of irredundance and zero forcing irredundance.
In light of Corollary 1.2, it is reasonable to consider replacing private forts in Definition 1.5 with private minimal forts. However, this modification does not align with the definition of irredundance: $v$ is a private neighbor of $u$ if $S\cap N[v]=\{u\}$, and this definition does not require that $S\cap N[x]=\{u\}$ implies that $N[v]\subseteq N[x]$. It also has an undesired consequence: For every graph $G$ and every vertex $v\in V(G)$, the set $\{v\}$ is an irredundant set. The same statement for $Z$-irredundance does not necessarily hold when restricting to private minimal forts. In particular, it does not hold for $Z$-irrelevant vertices (introduced in [6]), i.e., vertices that are not contained in any minimal zero forcing sets. By Proposition 1.7, a vertex is $Z$-irrelevant if and only if it is not contained in any minimal fort.
Remark 1.6.
Let $G$ be a graph and let $B$ be a minimal zero forcing set. Let $B^{\prime}=\{u\in B:u\mbox{ is in some minimal fort of }G\}$. For every fort $F$, $F$ contains a minimal fort $F^{\prime}$. Since $B$ is a zero forcing set, $\emptyset\neq B\cap F^{\prime}=B^{\prime}\cap F^{\prime}\subseteq B^{\prime}\cap F$, so $B^{\prime}$ is a zero forcing set. Since $B$ is minimal, $B=B^{\prime}$. So each element of $B$ is an element of some minimal fort of $G$.
Proposition 1.7.
Let $G$ be a graph and let $v\in V(G)$. Then $v$ is $\operatorname{Z}$-irrelevant if and only if $v$ is not in any minimal fort of $G$.
Proof.
If $v$ is not in any minimal fort, then $v$ is not in any minimal zero forcing set by Remark 1.6, so $v$ is $\operatorname{Z}$-irrelevant. Now assume $w$ is in a minimal fort $F_{0}$. Let the minimal forts of $G$ be denoted by $F_{0},F_{1},\dots,F_{k}$. Since each $F_{i}$ is minimal, $F_{0}\not\subseteq F_{i}$ for $i=1,\dots,k$. For each $i=1,\dots,k$, choose $v_{i}\in F_{i}\setminus F_{0}$ and define $\hat{B}=\{w\}\cup\{v_{i}:i=1,\dots,k\}$. Then $\hat{B}$ is a zero forcing set by Corollary 1.2, so $\hat{B}$ contains a minimal zero forcing set $B$. Since $v_{i}\neq F_{0}$, necessarily $w\in B$. ∎
@@ 4 Bounds
First, we assume that $c_{j-1}$ and $a_{j}$ cannot be in a private fort of the vertices in $V_{j}\cap S$, and show $|V_{j}\cap S|\leq 1$. Let $x\in\{b_{j},c_{j},d_{j}\}\cap S$ and let $F_{x}$ be a private fort of $x$. If $F_{x}\cap\{b_{j},d_{j}\}\neq\varnothing$, then $a_{j}$ must be adjacent to two vertices in the fort and $c_{j-1}\not\in F_{x}$, so $\{b_{j},d_{j}\}\subseteq F_{x}$. If $x=c_{j}$, then $b_{j}\in F_{x}$ or $b_{j}$ needs two neighbors in $F_{x}$ so $d_{j}\in F_{x}$, so $\{b_{j},c_{j},d_{j}\}\subseteq F_{x}$. Therefore $|V_{j}\cap S|\leq 1$. By symmetry, if $c_{j-1}$ and $a_{j}$ cannot be in a private fort of the vertices in $V_{j-1}\cap S$, then $|S\cap V_{j-1}|\leq 1$. Thus